OpenStudy (anonymous):

Can someone show me the sketch of x^2/9 – y^2/4 = 1 please.

jimthompson5910 (jim_thompson5910):

if you can, I recommend downloading geogebra it is a graphing calculator program that is very useful link: http://www.geogebra.org/cms/en/

jimthompson5910 (jim_thompson5910):

with this program, you can type in the exact equation above (you can copy and paste even) into the input bar and it will graph it for you

OpenStudy (anonymous):

wolf will do it too but you probably need to know that it is a hyperbola to start with and also know for example that the center is \((0,0)\) and some other facts

OpenStudy (anonymous):

form is \[\frac{(x-h)^2}{a^2}-\frac{(y-k)^2}{b^2}=1\] in your example \(h=0, k=0, a=3, b=2\)

jimthompson5910 (jim_thompson5910):

yes wolfram alpha is good too, but it's not as dynamic as it should be in my opinion maybe if you paid for their services, it would be better, idk

OpenStudy (anonymous):

okay, i see it now. i like having all of the information right in front of me! ill be sure to use this more often. thanks so much for the helpful site!!

OpenStudy (anonymous):

go with the first one